Question 1) A flea jumps by exerting a force of 1.22 10-5 N straight down on the ground. A breeze blowing on the flea parallel to the ground exerts a force of 1.12 10-6 N on the flea. Find the direction and magnitude of the acceleration of the flea if its mass is 6.0 10-7 kg. (Let us assume that Fwind points to the right. We will consider this to be the +x direction and vertical to be the +y direction.)

Question 2) Suppose two children push horizontally, but in exactly opposite directions, on a wagon containing a third child. The first child exerts a force of F1 = 75.0 N, the second a force of F2 = 91.0 N, friction is 12.0 N, and the mass of the third child plus wagon is 22.0 kg. What would the acceleration be if friction is 16.0 N?

Question 3) The 1800 kg tractor exerts a force of 1.90 104 N backward on the pavement, and the system experiences forces resisting motion that total 2400 N. If the acceleration is 0.150 m/s2, what is the mass of the airplane?
Answer #1

Q1

flea exerting a force down on the ground equal reaction force will be upward F = 1.22*10^-5 N

breeze parallel to the ground along + x direction is F2 = 1.12*10^-6 N

mass of flea is m = 6.0*10^-7 kg

writing the forces of

F1 = 0 N i +1.22*10^-5 N j

F2 = 1.12*10^-6 N i + 0 N j

net force is F = F1+F2 = (0 N i +1.22*10^-5 N j )+( 1.12*10^-6 N i + 0 N j)

F = 1.12*10^-6 N i + 1.22*10^-5 N j

the magnitude is F = sqrt((1.12*10^-6)^2+(1.22*10^-5)^2) N = 1.225130*10^-5 N

the direction is theta = arc tan ((1.22*10^-5)/(1.12*10^-6)) degrees = 84.75 degrees

the magnitude of the acceleration is a = F/m = (1.225130*10^-5)/(6*10^-7) m/s2 = 20.42 m/s2

Q2

F1 = 75 N , F2 = 91 N

the friction is f = 12 N

mass of third child and wagon is m = 22 kg

net force is (F1+f)-(F2+f) = m*a

(75+12)-(91-12)= 22*a

a = 0.3636 m/s2

if the friction force is 16 N thne

F1+f)-(F2+f) = m*a

(75+16)-(91-16)= 22*a

a = 0.7272 m/s2

Q3

net force is 19000 N - 2400 N = 16600 N

the acceleration is m = F/0.15 = 16600/0.15 = 110666.67 kg

so the mass of the plane is m = 110666.67 kg